OpenStudy (anonymous):

Multiplying rational expressions: 4x^2+20x/x^3+4x^2 * (x^2+8x+16)

OpenStudy (anonymous):

4(x+5)(x+4)/x

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

4x^2+20x/x^3+4x^2 * (x^2+8x+16) =4x(x+5)/x^2(x+4)*(x+4)^2 =4(x+5)/x(x+4)*(x+4)^2 =4(x+5)(x+4)/x

OpenStudy (anonymous):

Is that clear?

OpenStudy (anonymous):

Luci?

OpenStudy (anonymous):

so you factored the top of, correct? and got 4x(x+5)

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

i don't get how you got that, the factoring of the top i mean.

OpenStudy (anonymous):

or the bottom. :////

OpenStudy (anonymous):

wow. im a moron. i got it. sorry.

OpenStudy (anonymous):

and you do the same thign for the bottom one, right?

OpenStudy (anonymous):

Ya

OpenStudy (anonymous):

ahhhh.... got it:) sorry about that.

OpenStudy (anonymous):

Dat's k.